can you help me please movie tickets at a local theater cost $6 for adults and $2 dollars for kids under twelve.if 175 tickets were sold and that equals to $750 how many kid tickets were bought ?

Answer by mananth(16946)

Adults x tickets , Price = $6.00
children y tickets , Price = $2.00
Total 750
1.00 x + 1 y = 175 .............1
6 x + 2 y = 750 .............2
Eliminate y
multiply (1)by -2
Multiply (2) by 1
-2 x -2 y = -350
6 x + 2 y = 750
Add the two equations
4 x = 400
/ 4
x = 100
plug value of x in (1)
1 x + 1 y = 175
100 + 1 y = 175
1 y = 175 -100
1 y = 75
y = 75
Adults tickets 100
children tickets 75
m.ananth@hotmail.ca